# The Generation Phases of Flow Matching: a Denoising Perspective

### v1

@@ 5.2 Inpainting
We complement our experiments with a third metric, at the crossroads of denoising and generation, to confirm the correlation previously observed. To this end, we turn to an intermediate task: image inpainting. This task is both an inverse problem, a field in which denoisers are regularly used in a plug-and-play fashion (Venkatakrishnan et al., 2013; Meinhardt et al., 2017), and a generative challenge, as the model must synthesize large missing parts of the image. We evaluate our denoisers in this setting using the PnP-Flow algorithm (Martin et al., 2025), which incorporates FM models into plug-and-play frameworks. This method builds on the equivalence between denoisers and FM velocities, since in PnP-Flow the time-dependent denoiser is directly induced by the learned velocity field. Experimental details and visual examples are provided in Appendix E, Figure 9. As shown in Figure 3, the ranking of models on the inpainting task coincides with their ranking in terms of FID and PSNR. Now that it is clear that the considered models are uniformly good or bad across all three metrics, we investigate in more depth the factors that may cause this performance gap.
@@
6 The distinct influences of early and late times on generation
The above experiments revealed striking differences of performance between models trained with different losses; naturally raising the question: why are some of the considered models much worse at generating than the FM baseline and at which stage of the generative process do they fail? A first step towards addressing this question is to develop a more fine-grained understanding of the temporal structure underlying the generation process.
6.1 Perturbating artificially the generation process
We first implement controlled perturbations of the denoiser, applied at different selected time intervals during the generation process.
@@ 6.1 Perturbating artificially the generation process
    the (deterministic) perturbation direction $\delta$. We consider (a) checkerboard perturbations corresponding to alternated patches of +1+1 and 1-1, with different patch sizes; (b) positive (resp. negative) shift respectively referring to a constant perturbation of +1+1 (resp. 1-1); (c) “Residual” referring to the perturbation $\tilde{D}_{t}:=D_{t}+\sigma(t)(\mathrm{Id}-D_{t})$ which can be seen as a relaxed denoiser. Perturbations are displayed in Appendix F, Figure 10.
Figure 4: Influence of different perturbations at different generation phases on the FID. Two classes of perturbations emerge: noise-type perturbations (checkboard 4x4, residual) characterized by high FID, low pairwise distance and strongest impact in the last times and drift-type perturbations (pos./neg. shift, checkerboard 16x16) characterized by low FID, high pairwise distance and strongest impact in the early times.
Results are presented Figure 4. The left plot displays the average $\ell_{2}$ distance between $500$ generated samples and their corresponding baseline samples (i.e., generated with $D_{t}$ starting from the same noise instance $x_{0}$). This allows to measure how the injected noise deviates the ODE trajectory and affects the generated sample. The right plot displays the test FID-10k for each perturbation applied on each interval.
We observe that the $1\times 1$ checkerboard perturbation does not induce any significative change in the generated samples, meaning it is effectively corrected after the perturbed time interval (although, by design of the experiment, like all perturbations it degrades the PSNR by 10 %). As the size of the patches grows from $1\times 1$ to $16\times 16$, the generated images deviate further from the initial samples. Something surprising happens: although larger patches always induce higher drift in the distribution (as assessed by the pairwise distance), we observe that applying the $4\times 4$ kernel size at later times produces a large increase in FID, while for the other checkboard perturbations, they remain quite low. Regarding the constant perturbations, they strongly alter the generated images, producing washed-out outputs (see Appendix F, Figures 11 and 12) but yield only a small FID change. On the contrary, the residual perturbation leads to a strong increase of the FID, although the generated images are noisy versions, close to the initial ones in $\ell_{2}$ distance and visually.
@@
    First, despite all perturbations having, by design, the same impact on PSNR degradation, they do not affect FID in the same manner. A first remark is that it is possible to build denoisers with degraded denoising performance that still remain good generators. Second, the experiment shows a stronger sensitivity to noise-type perturbations (i.e. low patch-size checkerboard pattern or residual, both being Gaussian-like) than to drift-type ones (i.e. positive shift, negative shift, large patch-size checkerboard).
@@
    Second, drift-type perturbations, which induce global change to the full image, are most impactful when applied early, while noise-type perturbations, which act locally, have a stronger effect when applied in late time intervals. This aligns with previous studies on the effective receptive field of the velocity U-Net during generation: Kamb and Ganguli (2025, Fig. 4.a) and Niedoba et al. (2025, Fig. 3) show that it evolves from a large kernel that encompasses the full image (enabling the model to compute the average of the dataset at $t=0$, matching the closed-form target) to a local field (a few pixels) for the last time steps, corresponding to removal of very small noise.
@@
(a) Splitting trajectories in 1D flow matching between a uniform $p_{0}$ and a discrete target $\hat{p}_{1}$, composed of 3 data points. Red arrows represent the optimal velocity $\hat{v}^{\star}(x_{t},t)$.
@@
(b) Mean and standard deviation of the spectral norm $\|\nabla_{x}v(x_{t},t)\|_{2}$ computed on 1000 ODE trajectories on Cifar10 with Gaussian $p_{0}$.
@@
6.2 Regularity discrepancies between target and learned models
Previous experiments showed the influence of the early phase of the generative process, sensitive to drift-type perturbations, and late phase, sensitive to noise-type perturbations. This distinction of small and large times has already been observed in different contexts: Biroli et al. (2024) identify different temporal regimes in the generative process, evolving from trajectories that are indistinguishable to trajectories that all converge toward the training dataset, thereby revealing a phase transition between generalisation and memorisation. Bertrand et al. (2025) show that learned velocities differ the most from the loss minimizer $\hat{v}^{\star}$ (4) at small and large $t$, and that small $t$ seem to be more important for creating new images.
To deepen our understanding of the various phases, we consider a simpler setup with $p_{0}$ the uniform distribution on $1,1]^{d}$, and the discrete target distribution $\hat{p}_{1}$. In this setting, the optimal velocity field $\hat{v}^{\star}$ admits a simple closed-form expression (Bertrand et al., 2025, Prop. 1). It is defined on cones  [^4] For a Gaussian $p_{0}$, the optimal velocity $\hat{v}^{\star}_{t}$ is defined on the whole space: the regions not covered by cones in the uniform case correspond to regions of very low probability in the Gaussian setting. $C^{(i)}=\{x\in\mathbb{R}^{d}:\exists t\in[0,1],x_{0}\in[-1,1]^{d},\,x=(1-t)x_{0}+tx^{(i)}\}$. The optimal velocity $\hat{v}^{\star}(x_{t},t)$ simply points towards the mean of training points $x^{(i)}$ for all the indices $i$ such that $x_{t}\in C^{(i)}$ (Bertrand et al., 2025, Prop. 1). It follows that for $t$ close to 0, $\hat{v}^{\star}(x_{t},t)$ points towards the mean of the dataset, whereas in the later steps, $x_{t}$ belongs to a single cone and $\hat{v}^{\star}(x_{t},t)$ points towards the associated data point (Figure 5(a)). In between, there is a transition phase when the trajectories “split” into different cones.
This simple setting examplifies the two phases of the target velocity: during small times, it rapidly changes; then, after some threshold $\tau$, points $x$ only belong to a single cone $\mathcal{C}^{(i)}$ and the velocity, equal to $\frac{x^{(i)}-x}{1-t}$, varies very smoothly. We conjecture that approximating the closed-form is the hardest at such critical $\tau$ because of a high local Lipschitz constant of the target velocity field with respect to $x$, making it difficult to accurately capture the trajectory splitting dynamics.
To test this hypothesis numerically, on Figure 5(b), we estimate the local Lipschitz constant of the velocity in $x$ at time $t\in[0,1]$ by computing the spectral norm of the Jacobian $\nabla_{x}v(x_{t},t)$ along sampled ODE trajectories, using the power method, both for the closed-form $\hat{v}^{\star}$ and for velocities induced by our set of denoisers. This quantity reflects how strongly trajectories diverge locally. Once again, we identify distinct temporal behaviours. First, in an early regime ($t\in[0.1,0.2]$), the closed-form velocity exhibits a sharp Lipschitz peak, when trajectories split into cones, which confirms our intuition. Our trained models, by contrast, fail to reproduce this peak. The pronounced gap – at a phase already identified as critical for generalisation (Bertrand et al., 2025) – shows that networks learn a smoother velocity, unable to match the closed-form, and that this actually helps trajectories drift away from training samples thereby favoring generalisation. Second, in an intermediate regime ($t\in[0.3,0.8]$), the models maintain a relatively high Jacobian spectral norm, consistently above the closed-form, with the best-performing models showing the largest values in this range. This shows that maintaining a high Lipschitz constant during intermediate times is not problematic; instead, it may reflect the capacity of neural networks to represent complex transformations (Salmona et al., 2022).
@@ 6.3 Probing the temporal phases with ad-hoc denoisers
In Section 6.1, we investigated how to artificially perturb a pretrained denoiser, identifying two types of behaviors: early-time drifts and late-time FID degradation. We now ask whether these effects can be reproduced directly through training, by modifying either the loss function or the class of functions over which the loss is minimized, while remaining within the framework of our denoising toolkit.
@@
In order to explore the importance of matching the Lipschitz peak at $t=0.2$ of the closed form (see Figure 5(b)), we build a denoiser whose Jacobian spectral norm is softly penalized during training on early times (e.g. $[0.1,0.3]$). We fix as setup loss $\mathcal{L}_{\text{FM}}$, parametrization $\mathcal{C}_{\mathrm{I+NN}}$ and additional regularization denoted $\mathcal{R}_{\text{[0.1,0.3]}}$. Interestingly, this model matches the best performing model (standard FM) in FID with a Jacobian spectral norm that is twice the difference in interval $[0.1,0.3]$ (see Figures 14 and 15). This suggests that the early critical time identified for the closed-form is actually not the only decisive time when considering learned models. Conducting the mirror experiment, with regularization applied at late times, e.g. $\mathcal{R}_{\text{[0.5-0.8]}}$ applied on the interval $[0.5,0.8]$, leads to degraded generation performance. More precisely, while regularizing at early times produces models with similar FID than the FM baseline but that can generate visually different samples (see Figure 18), applying regularization at late times instead produces samples that are closer to the baseline models but with degraded FID, confirming the behaviors observed under artificial perturbations. We push further these experiments on regularized models in Section H.1.
@@
In the same spirit of dissecting which temporal regions of the trajectory matter most, we now train models with ad-hoc loss weightings that deliberately bias learning toward specific times. We build a new denoiser with weighting $w_{t}^{\mathrm{mid}}=\frac{1}{(0.5-t)^{2}}$, putting the emphasis on accurate denoising at $t=0.5$ whereas traditional weights focus on small and large $t$ (Kim et al., 2025, Fig. 2). We test and analyze other mid weightings in Section H.2. Figure 6, displaying the distance between generated samples starting from the same noise, shows that $w_{t}^{\text{mid}}$ induces substantial deviations from both FM/den models, while generating points that are roughly at the same distance of the dataset, and having an FID similar to $\mathcal{L}_{\mathrm{den}}$ (full evaluation is in Appendix C). This model strongly differs from the others (see also generated samples in Figures 13 and 18), an interesting fact as several works suggest, on the contrary, that all models, irrespective of architecture and optimization (Niedoba et al., 2025), and even subset of training data (Kadkhodaie et al., 2024) end up generating the same data. [Uncaptioned image] Figure 6: Left: average pairwise distance inter-models computed on $500$ samples, sharing same $x_{0}$ across models (CelebA-64). Right: average distance of samples to train set. $\mathcal{L}_{\mathrm{mid}}$ produces samples that differ from those of other models.
@@
    Similar FIDs can hide different generation behaviours. Our general denoiser framework makes it possible to explicitly build such models. By acting on the early/mid time of the generation process, we are able to change the samples generated.
@@ 7 Conclusion
Appendix A Details on the losses

### v2

@@ 4.2 Inpainting
We complement our experiments with a third metric, at the crossroads of denoising and generation, to confirm the correlation previously observed. To this end, we turn to an intermediate task: image inpainting. This task is both an inverse problem, a field in which denoisers are regularly used in a plug-and-play fashion (venkatakrishnan2013plug; meinhardt2017learning), and a generative challenge, as the model must synthesize large missing parts of the image. We evaluate our denoisers in this setting using the PnP-Flow algorithm (martin2024pnp), which incorporates FM models into plug-and-play frameworks. This method builds on the equivalence between denoisers and FM velocities, since in PnP-Flow the time-dependent denoiser is directly induced by the learned velocity field. Experimental details and visual examples are provided in Appendix F, Figure 10. As shown in Figure 3, the ranking of models on the inpainting task coincides with their ranking in terms of FID and PSNR. Now that it is clear that the considered models are uniformly good or bad across all three metrics, we investigate in more depth the factors that may cause this performance gap.
@@
5 Perturbating the generation process
The above experiments revealed striking differences of performance between models trained with different losses; naturally raising the question: why are some of the considered models much worse at generating than the FM baseline and at which stage of the generative process do they fail? A first step towards addressing this question is to develop a more fine-grained understanding of the temporal structure underlying the generation process. To do so, we implement controlled perturbations of the denoiser, applied at different selected time intervals during the generation process.
@@ 5 Perturbating the generation process
    the (deterministic) perturbation direction $\delta$. We evaluate both low- and high-frequency perturbation patterns. We consider (a) checkerboard perturbations corresponding to alternated patches of +1+1 and 1-1, with different patch sizes; (b) positive (resp. negative) shift respectively referring to a constant perturbation of +1+1 (resp. 1-1); (c) “Residual” referring to the perturbation $\tilde{D}_{t}:=D_{t}+\sigma(t)(\mathrm{Id}-D_{t})$ which can be seen as a relaxed denoiser. (d) low-pass–filtered Gaussian perturbations, corresponding to $\delta=h\star g_{0}$ where $g_{0}$ is a fixed Gaussian noise realization and $h$ is a low-pass filter with various cutoffs. Perturbations are displayed in Appendix G, Figures 12 and 13.
Figure 4: Influence of different perturbations at different generation phases on the FID (10K, test) on CelebA 128. Two classes of perturbations emerge: high-frequency perturbations (ckb. 4x4, 8x8) characterized by high FID, low pairwise distance and strongest impact in the last times and low-frequency perturbations (pos./neg. shift, ckb. 16x16-64x64) characterized by low FID, high pairwise distance and strongest impact in the early times.
Results are presented Figure 4 and in Appendix G. The left plot displays the average $\ell_{2}$ distance between $500$ generated samples and their corresponding baseline samples (i.e., generated with $D_{t}$ starting from the same noise instance $x_{0}$). This allows to measure how the injected noise deviates the ODE trajectory and affects the generated sample. The right plot displays the test FID-10k for each perturbation applied on each interval.
We observe that the $1\times 1$ checkerboard perturbation does not induce any significative change in the generated samples, meaning it is effectively corrected after the perturbed time interval (although, by design of the experiment, like all perturbations it degrades the PSNR by 10 %). As the size of the patches grows from $1\times 1$ to $16\times 16$, the generated images deviate further from the initial samples. Something surprising happens: although larger patches always induce higher drift in the distribution (as assessed by the pairwise distance), we observe that applying the $4\times 4$ kernel size at later times produces a large increase in FID, while for the other checkboard perturbations, they remain quite low. Regarding the constant perturbations, they strongly alter the generated images, producing washed-out outputs (see Appendix G, Figures 14, 18 and 15) but yield only a small FID change. On the contrary, the residual perturbation leads to a strong increase of the FID, although the generated images are noisy versions, close to the initial ones in $\ell_{2}$ distance and visually.
@@
    First, despite all perturbations having, by design, the same impact on PSNR degradation, they do not affect FID in the same manner. A first remark is that it is possible to build denoisers with degraded denoising performance that still remain good generators. Second, the experiment shows a stronger sensitivity to high-frequency perturbations (i.e. low patch-size checkerboard pattern or residual, both being Gaussian-like) than to low-frequency perturbations (i.e. positive shift, negative shift, large patch-size checkerboard).
@@
    Second, low-frequency perturbations, which induce a drift in the sampled distribution (i.e. a global change to the full image) are more impactful when applied early. In contrast, high-frequency perturbations, which produce noise-like local changes, have a stronger effect when applied in late time intervals. This aligns with previous studies on the effective receptive field of the velocity U-Net during generation: kamb2024analytic and niedoba2024towards show that it evolves from a large kernel that encompasses the full image (enabling the model to compute the average of the dataset at $t=0$, matching the closed-form target) to a local field (a few pixels) for the last time steps, corresponding to removal of very small noise.
No perturb.
Low-frequency High-frequency
Interval Neg. shift Ckb. $\times64$ Ckb. $\times8$ Ckb. $\times4$
$[0.0,0.3]$ Refer to caption Refer to caption Refer to caption Refer to caption
$[0.6,0.9]$ Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Effect of low- and high-frequency perturbations at early and late times for CelebA 128.
6 The distinct influences of early and late times on generation
Previous experiments showed the influence of the early phase of the generative process, sensitive to low-frequency perturbations with a drift effect on samples and late phase, sensitive to high-frequency perturbations with a noise effect on samples. This distinction of small and large times has already been observed in different contexts: biroli2024dynamical identify different temporal regimes in the generative process, evolving from trajectories that are indistinguishable to trajectories that all converge toward the training dataset, thereby revealing a phase transition between generalisation and memorisation. bertrand2025closed show that learned velocities differ the most from the loss minimizer $\hat{v}^{\star}$ (4) at small and large $t$, and that small $t$ seem to be more important for creating new images.
6.1 Regularity discrepancies between target and learned models
[…]
@@ 6.1 Regularity discrepancies between target and learned models
Figure 6: Splitting trajectories in 1D flow matching between a uniform $p_{0}$ and a discrete target $\hat{p}_{1}$, composed of 3 data points. Red arrows represent the optimal velocity $\hat{v}^{\star}(x_{t},t)$.
This setting examplifies distinct phases of the target closed-form velocity/denoiser plotted on Figure 7(a):
  • Very early phase. The closed form denoiser outputs the mean of the dataset: this is the best estimate at very high noise levels. Since the target denoiser is constant at $t=0$, its Jacobian norm is zero at these initial times.
  • Early phase. The closed-form velocity/denoiser rapidly changes, visible as a peak in the Jacobian norm. This corresponds to the splitting of ODE trajectories.
  • Intermediate phase. After some time threshold $\tau$, points $x_{t}$ only belong to a single cone $\mathcal{C}^{(i)}$ and the velocity, equal to $\frac{x^{(i)}-x}{1-t}$, varies very smoothly. Equivalently, the denoiser’s output is constant equal to $x^{(i)}$ and its Jacobian norm vanishes.
  • Late phase. Due to the factor 1/(1t)1/(1-t) in the velocity formula, the Jacobian norm of the target velocity explodes while the target denoiser’s one remains null.
We conjecture that approximating the closed-form is the hardest at such critical $\tau$ because of a high local Lipschitz constant of the target velocity field with respect to $x$, making it difficult to accurately capture the trajectory splitting dynamics. We also conjecture that it is desirable to maintain a higher Lipschitz constant that the target velocity / denoiser in the intermediate phase.
To test this hypothesis numerically, on Figure 7(b), we estimate the local Lipschitz constant of the velocity in $x$ at time $t\in[0,1]$ by computing the spectral norm of the Jacobian $\nabla_{x}v(x_{t},t)$ along sampled ODE trajectories, using the power method, both for the closed-form $\hat{v}^{\star}$ and for velocities induced by our set of denoisers. This quantity reflects how strongly trajectories diverge locally.
Once again, we identify distinct temporal behaviours. First, in the early phase ($t\in[0.1,0.3]$), the closed-form velocity exhibits a sharp Lipschitz peak, while our trained models, by contrast, fail to reproduce this peak; which confirms our intuition. The pronounced gap - at a phase already identified as critical for generalisation (bertrand2025closed) - shows that networks learn a smoother velocity, unable to match the closed-form, and that this actually helps trajectories drift away from training samples thereby favoring generalisation. Second, in the intermediate regime ($t\in[0.3,0.8]$), the models maintain a relatively high Jacobian spectral norm, consistently above the closed-form, with the best-performing models showing the largest values in this range. This shows that maintaining a high Lipschitz constant during intermediate times is not problematic; instead, it may reflect the capacity of neural networks to represent complex transformations (salmona2022can).
@@
(a) Mean of the spectral norm $\|\nabla_{x}D(x_{t},t)\|_{2}$ computed on 1000 ODE trajectories on Cifar10 with Gaussian $p_{0}$. Colored regions indicate the temporal phases defined by the closed-form dynamics.
(b) Mean of the spectral norm $\|\nabla_{x}v(x_{t},t)\|_{2}$ computed on 1000 ODE trajectories on Cifar10 with Gaussian $p_{0}$.
@@ 6.2 Probing the temporal phases with ad-hoc denoisers
In Section 5, we investigated how to artificially perturb a pretrained denoiser, identifying two types of behaviors: early-time drifts and late-time FID degradation. We now ask whether these effects can be reproduced directly through training, by modifying either the loss function or the class of functions over which the loss is minimized, while remaining within the framework of our denoising toolkit.
@@
In order to explore the importance of matching the Lipschitz peak at $t=0.2$ of the closed form (see Figure 7(b)), we build a denoiser whose Jacobian spectral norm is softly penalized during training on early times (e.g. $[0.1,0.3]$). We fix as setup loss $\mathcal{L}_{\text{FM}}$, parametrization $\mathcal{C}_{\mathrm{I+NN}}$ and additional regularization denoted $\mathcal{R}_{\text{[0.1,0.3]}}$. Interestingly, this model matches (almost) the best performing model (standard FM) in FID with a Jacobian spectral norm that is twice the difference in interval $[0.1,0.3]$ (see Figures 19 and 20). Conducting the mirror experiment, with regularization applied at intermediate times, e.g. $\mathcal{R}_{\text{[0.4-0.8]}}$ applied on the interval $[0.4,0.8]$, leads to degraded generation performance. More precisely, while regularizing at early times produces a model with similar FID than the FM baseline but that can generate visually different samples (see Figure 23), applying regularization at intermediate times instead produces samples closer to the baseline models but with degraded FID, confirming the behaviors observed under artificial perturbations. We push further these experiments on regularized models in Section H.1.
In the same spirit of dissecting which temporal regions of the trajectory matter most, we now train models with ad-hoc loss weightings that deliberately bias learning toward specific times. We build a new denoiser with weighting $w_{t}^{\mathrm{mid}}=\frac{1}{(0.5-t)^{2}}$, putting the emphasis on accurate denoising at $t=0.5$ whereas traditional weights focus on small and large $t$ (kim2024simple, Fig. 2). We test and analyze other mid weightings in Section H.2.
@@
Figure 8, displaying the distance between generated samples starting from the same noise, shows that $w_{t}^{\text{mid}}$ induces substantial deviations from both FM/den models, while generating points that are roughly at the same distance of the dataset, and having an FID similar to $\mathcal{L}_{\mathrm{den}}$ (full evaluation is in Appendix D). This model strongly differs from the others (see also generated samples in Figure 24, Figure 25), an interesting fact as several works suggest, on the contrary, that all models, irrespective of architecture and optimization (niedoba2024towards; zhang2023emergence), and even subset of training data (kadkhodaie2024generalisation) end up generating the same data.
Figure 8: Left: average pairwise distance inter-models computed on $500$ samples, sharing same $x_{0}$ across models (CelebA-64). Right: average distance of samples to train set. $\mathcal{L}_{\mathrm{mid}}$ produces samples that differ from those of other models.
@@
    By acting on the early/mid time of the generation process, we are able to change the samples generated while maintaining reasonably low FIDs. Our general denoiser framework makes it possible to explicitly build such models.
@@ 7 Conclusion
Appendix A Related works
Weighting strategies in denoising score matching.
We now discuss related works within the diffusion framework, which is known to be equivalent to flow matching when the source distribution $p_{0}$ is Gaussian [gao2025diffusion]. For simplicity, we consider the variance-preserving diffusion process where $x_{t}$ is obtained from a clean sample $x_{0}$ and a Gaussian noise $\varepsilon\sim\mathcal{N}(0,I_{d})$ as $x_{t}:=\alpha_{t}x_{0}+\sigma_{t}\varepsilon$ with $\alpha_{t}^{2}+\sigma_{t}^{2}=1$. Note that we use the standard diffusion notations in this section only, meaning that the sampling process evolves from $t=T$ (noise) to $t=0$ (clean image), as opposed to the flow matching convention where $t$ evolves from 0 to $1$. Regarding parametrization, most diffusion implementations train a network to predict the noise $\varepsilon$. Some papers [salimans2022progressive, hang2023efficient] also study $v$-prediction (i.e. $\mathcal{C}_{\mathrm{I+NN}}$) or $x$-prediction (i.e. $\mathcal{C}_{\mathrm{NN}}$).
Regarding weighting, most diffusion papers follow the formulation from Ho2020 which uses an unweighted loss in the $\varepsilon$-prediction, i.e. $\mathbb{E}_{t,x_{0},\varepsilon}[\|\varepsilon-\varepsilon^{\theta}(x_{t},t)\|^{2}]$, corresponding in the denoising framework to the weighting ${\alpha_{t}^{2}}/{\sigma_{t}^{2}}$, equal to the signal-to-noise ratio. A line of work is devoted to crafting loss weightings. salimans2022progressive propose the weighting $\max(\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}},1)$ while yu2024unmasking use weighting $\frac{\alpha_{t}}{\sigma_{t}}$. Both approaches assign greater importance to large noise levels, arguing that these correspond to more difficult denoising tasks and play a crucial role in error propagation during sampling. Interpreting diffusion training as multitask optimization, hang2023efficient observe conflict between different timesteps objectives and suggest weighting $\min(\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}},\gamma)$ to avoid putting too much weight on the low noise levels, identified as an easy denoising task. In contrast, choi2022perception introduce the P2 weighting that puts more weight on the intermediate times, hypothesizing that perceptual features emerge during this content phase, as opposed to the early coarse phase or the late cleanup phase, where little noise remains in the image.
Rather than hypothesizing which timesteps are important or directly relating the weighting strategy to FID performance, we introduce an intermediate test metric–the PSNR evaluated at each timestep–which enables a more fine-grained analysis of how different timesteps contribute to generation quality. Similar to choi2022perception, we find that effective denoising at intermediate times is crucial for high-quality generation.
Unifying perspectives on loss weightings.
kingma2023understanding provide a unifying view on a wide range of loss weightings (including FM and the previous mentioned ones) by rewriting them as differently weighted ELBO formulations. kumar2025loss also lay down the different losses (noise prediction, score prediction, etc) and the loss weighting they induce.
Linking diffusion and denoising.
Beyond standard diffusion, delbracio2023inversion study a broader “degradation-to-clean” setting (which is not limited to Gaussian denoising) and pick a standard denoising loss (i.e. weighting $w_{t}^{\mathrm{den}}=1$, parametrization $\mathcal{C}_{\mathrm{NN}}$). For diffusion, they report good results on CelebA ($64\times 64$), yet below state-of-the-art levels, which is consistent with our observations. leclaire2025backward also provide a synthetic overview of the connections between diffusion models and classical additive Gaussian denoising, interpreting diffusion as an iterative “Noising-Relaxed Denoising” process to better understand the role of noise schedules.
Analysis of the generation phases.
One line of research focuses directly on the closed-form, either theoretically [biroli2024dynamical] or empirically [bertrand2025closed]. Both works highlight a critical time, at early timesteps, beyond which the closed-form points towards a single training example, while the trained models begin to deviate from it.
Another group of works seeks to understand the reasons for generalisation by characterizing the velocities or scores that are actually learned. kadkhodaie2024generalisation, niedoba2024towards, kamb2024analytic analyze the effective receptive field of trained models and show that it evolves from global (at high noise levels) to local (at low noise levels). niedoba2024towards, kamb2024analytic both argue that what is actually learned by the models is a patch-wise version of the closed-form, with patch size determined by the model’s receptive field: they demonstrate that this new formulation can predict in certain cases, without training, the samples learned by the model.
Beyond works focusing on the closed-form and its approximation, sclocchi2025phase connect the diffusion generative process to the learning of high- and low-level features: they specifically show that there exists a critical time at which image class is determined, while low-level features evolve smoothly throughout the generation process.
Finally, some works explain generalisation by imperfect or early-stopped optimization [wu2025taking, bonnaire2025diffusion, favero2025bigger].
Appendix B Details on the losses